# Compatible $E$-differential forms on Lie algebroids over (pre-)multisymplectic manifolds

### v3

@@ 6 Q-manifold description
$\left.+(-1)^{n}\frac{1}{n!}(\partial_{i}J_{a_{1}\ldots a_{n}}-\rho^{j_{1}}_{a_{1}}\ldots\rho^{j_{n}}_{a_{n}}H_{ij_{1}\ldots j_{n}})q^{a_{1}}\ldots q^{a_{n}}\right]\frac{\partial}{\partial z_{i}},$ (77)
@@
$H=\frac{1}{(n+1)!}H_{i_{1}\ldots i_{n+1}}(x)\mathrm{d}x^{i_{1}}\wedge\ldots\wedge\mathrm{d}x^{i_{n+1}},$ (78)
is a closed $n+1$-form. For a $1$-form $\alpha=\alpha_{i}(x)\mathrm{d}x^{i}$, $({\omega}_{grad}^{\flat})^{-1}:T^{*}[n-1]E[1]\rightarrow TE[1]$ is the map induced from the symplectic form ${\omega}_{grad}$.
$({\omega}_{grad}^{\flat})^{-1}\alpha=\alpha_{i}(x)\frac{\partial}{\partial z_{i}}.$ (79)
Proposition 6.3
Let $E$ be a Lie algebroid and $H$ is a pre-$n$-plectic form. The vector field of degree $+1$, ${Q_{\mathrm{comp}}}$ in (77) satisfies ${Q_{\mathrm{comp}}}^{2}=0$ if and only if $J$ is a compatible $E$-$n$-form.
@@
We compute the square of ${Q_{\mathrm{comp}}}$. Using equations $Q_{LieJ}^{2}=0$ and ${\{{{\Theta_{Lie}},{\iota_{\rho(q)}^{n}H}}\}}={\{{{{\underline{J}}{}},{\iota_{\rho(q)}^{n}H}}\}}=0$, etc., we obtain that ${Q_{\mathrm{comp}}}^{2}$ is proportional to
${\{{{\Theta_{Lie}},{{\underline{J}}{}}}\}}+\iota_{\rho(q)}^{n+1}H,$ (80)
which must be zero. This equation is equivalent to the condition of a compatible $E$-$n$-form (11). ∎
Example 6.1
Let $\pi=\frac{1}{2}\pi^{ij}(x)\partial_{i}\wedge\partial_{j}\in\Gamma(\wedge^{2}TM)$ be a bivector field and $H=\frac{1}{3!}H_{ijk}(x)\mathrm{d}x^{i}\wedge\mathrm{d}x^{j}\wedge\mathrm{d}x^{k}\in\Omega^{3}(M)$ be a closed $3$-form. We take $n=2$ and consider an N-manifold $T^{*}[1]T[1]M$. Local coordinates are $(x^{i},q_{i},z_{i},y^{i})$ of degree $(0,1,1,0)$. Now, the homological vector field $Q$ is given by
$Q$ $=\pi^{ij}(x)q_{i}\frac{\partial}{\partial x^{j}}-\frac{1}{2}\partial_{i}\pi^{jk}(x)q_{j}q_{k}\frac{\partial}{\partial q_{i}}-\left(\pi^{ij}z_{j}+\partial_{k}\pi^{ij}(x)q_{j}y^{k}-\pi^{ij}q_{j}\right)\frac{\partial}{\partial y_{i}}$
$+\left[\partial_{i}\pi^{jk}z_{j}q_{k}+\frac{1}{2}\partial_{i}\partial_{j}\pi^{kl}q_{k}q_{l}y_{j}-\frac{1}{2}(\partial_{i}\pi^{jk}-\pi^{jl}\pi^{km}H_{ilm})q_{j}q_{k}\right]\frac{\partial}{\partial z_{i}},$ (81)
@@
There does not necessarily exist a homological function $\Theta$ satisfying ${Q_{\mathrm{comp}}}={\{{{\Theta},{-}}\}}$ for the homological function in (77), if $H$ is nonzero. A compatible $E$-$n$-form can not be constructed as a QP-structure on $T^{*}[n-1]E[1]$.****** Since a homological function for a compatible $E$-$n$-form does not exist, we cannot construct a topological sigma model by using the so called AKSZ construction [2].
@@ 7 Realization by QP-manifold and twist
In the Q-manifold description in the previous section, a compatible condition is not formulated as a QP-manifold, i.e., there exists no homological function $\Theta$ such that $\iota_{Q}\omega=-\delta\Theta$ for the homological vector field $Q$ in Equation (77) if $H\neq 0$. We consider another realization in graded geometry as a Lagrangian Q-submanifold of a QP-manifold.
For it, we consider a shifted cotangent bundle of the graded bundle in Section 6. Take a shifted double cotangent bundle $T^{*}[n]{\mathcal{M}}$ of ${\mathcal{M}}=T^{*}[n-1]E[1]$ in the previous section 6. We consider a QP-structure on $T^{*}[n]{\mathcal{M}}$, i.e., we take the canonical graded symplectic form ${\widetilde{\omega}}_{grad}$ and ${\widetilde{Q}}$ such that ${\mathcal{L}}_{{\widetilde{Q}}}{\widetilde{\omega}}_{grad}=0$.
@@ 7.1 Twisting of homological functions and twisted QP-manifolds
Let ${pr}:T^{*}[n]{\mathcal{M}}\rightarrow{\mathcal{M}}$ be a natural projection map. ${\mathcal{M}}$ is a trivial Lagrangian graded submanifold since $\{{pr}^{*}f,{pr}^{*}g\}=0$ for all $f,g\in C^{\infty}({\mathcal{M}})$ and ${\mathcal{M}}$ is of half dimensions of $T^{*}[n]{\mathcal{M}}$.
If $(T^{*}[n]{\mathcal{M}},{\widetilde{\omega}}_{grad},{\widetilde{Q}})$ is a QP-manifold of degree $n$, an induced graded Poisson structure of degree $-n$ is defined on ${\mathcal{M}}$ as follows. A bilinear graded bracket of degree $-n+1$ is defined on ${\mathcal{M}}$ by a derived bracket,
$\{f,g\}_{{\mathcal{M}}}$ $:={pr}_{*}\{\{{pr}^{*}f,{\widetilde{\Theta}}\},{pr}^{*}g\},$ (82)
@@
$\{f,g\}_{{\mathcal{M}}}$ $=-(-1)^{(|f|-n+1)(|g|-n+1)}\{g,f\}_{{\mathcal{M}}},$ (83)
$\{f,gh\}_{{\mathcal{M}}}$ =$=\{f,g\}_{{\mathcal{M}}}h+(-1)^{(|f|-n+1)|g|}g\{f,h\}_{{\mathcal{M}}},$ (84)
$\{f,\{g,h\}_{{\mathcal{M}}}\}_{{\mathcal{M}}}$ =$=\{\{f,g\}_{{\mathcal{M}}},h\}_{{\mathcal{M}}}+(-1)^{(|f|-n+1)(|g|-n+1)}\{g,\{f,h\}_{{\mathcal{M}}}\}_{{\mathcal{M}}},$ (85)
@@
Definition 7.1
Let $\phi\in C^{\infty}(T^{*}[n]{\mathcal{M}})$. Then the twist $e^{\mathrm{ad}\phi}$ of any function $\alpha\in C^{\infty}({\mathcal{M}})$ by $\phi$ is defined by [37]
$e^{\mathrm{ad}\phi}\alpha:=\alpha+\{\alpha,\phi\}+\frac{1}{2}\{\{\alpha,\phi\},\phi\}+\ldots+\frac{1}{n!}\{\ldots\{\{\alpha,\phi\},\phi\},\ldots,\phi\}+\ldots.$ (86)
If a function $\phi\in C^{\infty}(T^{*}[n]{\mathcal{M}})$ is of degree $n$, twist preserves degree and gives a Poisson map $\alpha\mapsto e^{\mathrm{ad}\phi}\alpha$ satisfying
$\{e^{\mathrm{ad}\phi}\alpha,e^{\mathrm{ad}\phi}\beta\}=e^{\mathrm{ad}\phi}\{\alpha,\beta\},$ (87)
@@
Let ${\mathcal{E}}_{\phi}:=\{e^{\mathrm{ad}\phi}{pr}^{*}f|f\in C^{\infty}({\mathcal{M}})\}$ be a space of functions twisted by a degree $n$ function $\phi\in C^{\infty}(T^{*}[n]{\mathcal{M}})$. In general, ${\mathcal{E}}_{\phi}$ is identified as a space of functions on some graded manifold ${\mathcal{L}}_{\phi}\subset T^{*}[n]{\mathcal{M}}$, ${\mathcal{E}}_{\phi}=C^{\infty}({\mathcal{L}}_{\phi})$, by applying a Serre-Swan type theorem on a graded manifold.
Proposition 7.2
@@
It is confirmed using the formula $\{e^{\mathrm{ad}\phi}{pr}^{*}f,e^{\mathrm{ad}\phi}{pr}^{*}g\}=e^{\mathrm{ad}\phi}\{{pr}^{*}f,{pr}^{*}g\}=0$ and the dimension of ${\mathcal{L}}_{\phi}$ is the same as ${\mathcal{M}}$.
By definition, for $\alpha,\beta\in C^{\infty}({\mathcal{L}}_{-\phi})$, there exist $f,g\in C^{\infty}({\mathcal{M}})$ such that $\alpha=e^{-\mathrm{ad}\phi}{pr}^{*}f$ and $\beta=e^{-\mathrm{ad}\phi}{pr}^{*}g$. The derived bracket of $\alpha$ and $\beta$ is equal to
$\{\{e^{-\mathrm{ad}\phi}{pr}^{*}f,{\widetilde{\Theta}}\},e^{-\mathrm{ad}\phi}{pr}^{*}g\}$ $:=e^{-\mathrm{ad}\phi}\{\{{pr}^{*}f,e^{\mathrm{ad}\phi}{\widetilde{\Theta}}\},{pr}^{*}g\}.$ (88)
@@
$\{\alpha,\beta\}_{{\mathcal{L}}_{-\phi}}$ $:={pr}_{*}e^{-\mathrm{ad}\phi}\{\{{pr}^{*}f,e^{\mathrm{ad}\phi}{\widetilde{\Theta}}\},{pr}^{*}g\}.$ (89)
Since it is equivalent to ${pr}_{*}\{\{e^{-\mathrm{ad}\phi}{pr}^{*}f,{\widetilde{\Theta}}\},e^{-\mathrm{ad}\phi}{pr}^{*}g\}$, twisting by $\phi$ gives a graded Poisson bracket on ${\mathcal{L}}_{-\phi}$ induced by ${\widetilde{\Theta}}$ equivalent to a graded Poisson bracket on ${\mathcal{M}}$ induced by $e^{\mathrm{ad}\phi}{\widetilde{\Theta}}$. We can prove that {,}$\{-,-\}_{{\mathcal{L}}_{-\phi}}$ is a graded Poisson bracket of degree $-n+1$ similar to the proof for the bracket (82).
Thus, in the preceding discussion, we analyze the twisted homological function $e^{\mathrm{ad}\phi}{\widetilde{\Theta}}$ on ${\mathcal{M}}$ instead of a homological function on ${\mathcal{L}}_{-\phi}$.
The twisting function $\phi$ is an almost homological function. By (88) and (89), the graded Poisson bracket is defined by the derived bracket $\{\{-,{\widetilde{\Theta}}\},-\}$. On the other hand, we obtain
${pr}_{*}e^{\mathrm{ad}\phi}{\widetilde{\Theta}}$ $={pr}_{*}\ \left({\widetilde{\Theta}}+\{{\widetilde{\Theta}},\phi\}+\frac{1}{2}\{\{{\widetilde{\Theta}},\phi\},\phi\}+\ldots+\frac{1}{n!}\underbrace{\{\ldots\{\{{\widetilde{\Theta}},\phi\},\phi\},\ldots,\phi\}}_{n\,\text{brackets}}+\ldots\right),$ (90)
The function $\phi$ is of degree $n$, and from Equation (82), the third term of ${pr}_{*}\,e^{\mathrm{ad}\phi}\Theta$ gives
$\{\phi,\phi\}_{{\mathcal{M}}}=-{pr}_{*}\ \{\{{\widetilde{\Theta}},\phi\},\phi\}.$ (91)
If Equation (91) is zero, $\phi$ is a homological function on ${\mathcal{M}}$ and (,{,}$({\mathcal{M}},\{-,-\}_{{\mathcal{M}}},\phi)$ is a QP-manifold of degree $n-1$. However, Equation (91) is not necessarily zero.
Inspired by the above analysis and Equation (90), we define a twisted QP-structure. We consider the following class of generalizations of a QP-manifold which has good properties but it is not a QP.
Definition 7.3
@@
${pr}_{*}e^{\mathrm{ad}\phi}\Theta$ $=0,$ (92)
(,{,}$({\mathcal{M}},\{-,-\}_{{\mathcal{M}}},\phi)$ is called a twisted QP-manifold, where ${pr}:T^{*}[n]{\mathcal{M}}\rightarrow{\mathcal{M}}$ is a natural projection.
A function $\phi$ satisfying the condition (92) is also called a a Poisson function or a canonical function. Obviously a QP-manifold is a twisted QP-manifold. A twisted QP-manifold has some applications to physical theories such as the reduction of the AKSZ sigma models and the supergeometric construction of current algebras.[28, 29, 3]

### v4

@@ 6 Q-manifold description
$\left.+(-1)^{n}\frac{1}{n!}\bigl{(}\partial_{i}J_{a_{1}\dots a_{n}}-\rho^{j_{1}}_{a_{1}}\cdots\rho^{j_{n}}_{a_{n}}H_{ij_{1}\dots j_{n}}\bigr{)}q^{a_{1}}\cdots q^{a_{n}}\right]\frac{\partial}{\partial z_{i}},$ (6.3)
@@
$H=\frac{1}{(n+1)!}H_{i_{1}\dots i_{n+1}}(x)\mathrm{d}x^{i_{1}}\wedge\cdots\wedge\mathrm{d}x^{i_{n+1}},$
is a closed $(n+1)$-form. For a $1$-form
$\alpha=\alpha_{i}(x)\mathrm{d}x^{i},\qquad\bigl{(}{\omega}_{\rm grad}^{\flat}\bigr{)}^{-1}\colon\ T^{*}[n-1]E[1]\rightarrow TE[1]$
is the map induced from the symplectic form ${\omega}_{\rm grad}$.
$\bigl{(}{\omega}_{\rm grad}^{\flat}\bigr{)}^{-1}\alpha=\alpha_{i}(x)\frac{\partial}{\partial z_{i}}.$
Proposition 6.3.
Let $E$ be a Lie algebroid and $H$ is a pre-$n$-plectic form. The vector field of degree $+1$, ${Q_{\mathrm{comp}}}$ in (6.3) satisfies ${Q_{\mathrm{comp}}}^{2}=0$ if and only if $J$ is a compatible $E$-$n$-form.
@@
We compute the square of ${Q_{\mathrm{comp}}}$. Using equations $Q_{{\rm Lie}\,J}^{2}=0$ and
$\big{\{}\Theta_{\rm Lie},\iota_{\rho(q)}^{n}H\big{\}}=\big{\{}{\underline{J}}{},\iota_{\rho(q)}^{n}H\big{\}}=0,$
etc., we obtain that ${Q_{\mathrm{comp}}}^{2}$ is proportional to ${\{{{\Theta_{\rm Lie}},{{\underline{J}}{}}}\}}+\iota_{\rho(q)}^{n+1}H$, which must be zero. This equation is equivalent to the condition of a compatible $E$-$n$-form (3.1). ∎
Example 6.4.
Let $\pi=\frac{1}{2}\pi^{ij}(x)\partial_{i}\wedge\partial_{j}\in\Gamma\bigl{(}\wedge^{2}TM\bigr{)}$ be a bivector field and $H=\frac{1}{3!}H_{ijk}(x)\mathrm{d}x^{i}\wedge\mathrm{d}x^{j}\wedge\mathrm{d}x^{k}\in\Omega^{3}(M)$ be a closed $3$-form. We take $n=2$ and consider an N-manifold $T^{*}[1]T[1]M$. Local coordinates are $\bigl{(}x^{i},q_{i},z_{i},y^{i}\bigr{)}$ of degree $(0,1,1,0)$. Now, the homological vector field $Q$ is given by
$Q={}$ $\pi^{ij}(x)q_{i}\frac{\partial}{\partial x^{j}}-\frac{1}{2}\partial_{i}\pi^{jk}(x)q_{j}q_{k}\frac{\partial}{\partial q_{i}}-\bigl{(}\pi^{ij}z_{j}+\partial_{k}\pi^{ij}(x)q_{j}y^{k}-\pi^{ij}q_{j}\bigr{)}\frac{\partial}{\partial y_{i}}$
$+\left[\partial_{i}\pi^{jk}z_{j}q_{k}+\frac{1}{2}\partial_{i}\partial_{j}\pi^{kl}q_{k}q_{l}y_{j}-\frac{1}{2}\bigl{(}\partial_{i}\pi^{jk}-\pi^{jl}\pi^{km}H_{ilm}\bigr{)}q_{j}q_{k}\right]\frac{\partial}{\partial z_{i}},$
@@
There does not necessarily exist a homological function $\Theta$ satisfying ${Q_{\mathrm{comp}}}={\{{{\Theta},{-}}\}}$ for the homological function in (6.3), if $H$ is nonzero. A compatible $E$-$n$-form can not be constructed as a QP-structure on $T^{*}[n-1]E[1]$. [^6] Since a homological function for a compatible $E$-$n$-form does not exist, we cannot construct a topological sigma model by using the so called AKSZ construction [2].
@@ 7 Realization by QP-manifold and twist
In the Q-manifold description in the previous section, a compatible condition is not formulated as a QP-manifold, i.e., there exists no homological function $\Theta$ such that $\iota_{Q}\omega=-\delta\Theta$ for the homological vector field $Q$ in equation (6.3) if $H\neq 0$. We consider another realization in graded geometry as a Lagrangian Q-submanifold of a QP-manifold.
For it, we consider a shifted cotangent bundle of the graded bundle in Section 6. Take a shifted double cotangent bundle $T^{*}[n]{\mathcal{M}}$ of ${\mathcal{M}}=T^{*}[n-1]E[1]$ in the previous Section 6. We consider a QP-structure on $T^{*}[n]{\mathcal{M}}$, i.e., we take the canonical graded symplectic form ${\widetilde{\omega}}_{\rm grad}$ and ${\widetilde{Q}}$ such that ${\mathcal{L}}_{{\widetilde{Q}}}{\widetilde{\omega}}_{\rm grad}=0$.
@@ 7.1 Twisting of homological functions and twisted QP-manifolds
Let $\operatorname{pr}\colon T^{*}[n]{\mathcal{M}}\rightarrow{\mathcal{M}}$ be a natural projection map. ${\mathcal{M}}$ is a trivial Lagrangian graded submanifold since $\{\operatorname{pr}^{*}f,\operatorname{pr}^{*}g\}=0$ for all $f,g\in C^{\infty}({\mathcal{M}})$ and ${\mathcal{M}}$ is of half dimensions of $T^{*}[n]{\mathcal{M}}$.
If $\bigl{(}T^{*}[n]{\mathcal{M}},{\widetilde{\omega}}_{\rm grad},{\widetilde{Q}}\bigr{)}$ is a QP-manifold of degree $n$, an induced graded Poisson structure of degree $-n$ is defined on ${\mathcal{M}}$ as follows. A bilinear graded bracket of degree $-n+1$ is defined on ${\mathcal{M}}$ by a derived bracket,
$\{f,g\}_{{\mathcal{M}}}$ $:=\operatorname{pr}_{*}\big{\{}\big{\{}\operatorname{pr}^{*}f,{\widetilde{\Theta}}\big{\}},\operatorname{pr}^{*}g\big{\}},$ (7.1)
@@
$\{f,g\}_{{\mathcal{M}}}=-(-1)^{(|f|-n+1)(|g|-n+1)}\{g,f\}_{{\mathcal{M}}},$
$\{f,gh\}_{{\mathcal{M}}}=\{f,g\}_{{\mathcal{M}}}h+(-1)^{(|f|-n+1)|g|}g\{f,h\}_{{\mathcal{M}}},$
$\{f,\{g,h\}_{{\mathcal{M}}}\}_{{\mathcal{M}}}=\{\{f,g\}_{{\mathcal{M}}},h\}_{{\mathcal{M}}}+(-1)^{(|f|-n+1)(|g|-n+1)}\{g,\{f,h\}_{{\mathcal{M}}}\}_{{\mathcal{M}}}$
@@
Definition 7.1.
Let $\phi\in C^{\infty}(T^{*}[n]{\mathcal{M}})$. Then the twist ${\rm e}^{\operatorname{ad}\phi}$ of any function $\alpha\in C^{\infty}({\mathcal{M}})$ by $\phi$ is defined by [37]
${\rm e}^{\operatorname{ad}\phi}\alpha:=\alpha+\{\alpha,\phi\}+\frac{1}{2}\{\{\alpha,\phi\},\phi\}+\cdots+\frac{1}{n!}\{\dots\{\{\alpha,\phi\},\phi\},\dots,\phi\}+\cdots.$
If a function $\phi\in C^{\infty}(T^{*}[n]{\mathcal{M}})$ is of degree $n$, twist preserves degree and gives a Poisson map ${\alpha\mapsto{\rm e}^{\operatorname{ad}\phi}\alpha}$ satisfying
$\big{\{}{\rm e}^{\operatorname{ad}\phi}\alpha,{\rm e}^{\operatorname{ad}\phi}\beta\big{\}}={\rm e}^{\operatorname{ad}\phi}\{\alpha,\beta\}$
@@
Let ${\mathcal{E}}_{\phi}:=\big{\{}{\rm e}^{\operatorname{ad}\phi}\operatorname{pr}^{*}f\mid f\in C^{\infty}({\mathcal{M}})\big{\}}$ be a space of functions twisted by a degree $n$ function $\phi\in C^{\infty}(T^{*}[n]{\mathcal{M}})$. In general, ${\mathcal{E}}_{\phi}$ is identified as a space of functions on some graded manifold ${\mathcal{L}}_{\phi}\subset T^{*}[n]{\mathcal{M}}$, ${\mathcal{E}}_{\phi}=C^{\infty}({\mathcal{L}}_{\phi})$, by applying a Serre–Swan type theorem on a graded manifold.
Proposition 7.2.
@@
It is confirmed using the formula $\big{\{}{\rm e}^{\operatorname{ad}\phi}\operatorname{pr}^{*}f,{\rm e}^{\operatorname{ad}\phi}\operatorname{pr}^{*}g\big{\}}={\rm e}^{\operatorname{ad}\phi}\{\operatorname{pr}^{*}f,\operatorname{pr}^{*}g\}=0$ and the dimension of ${\mathcal{L}}_{\phi}$ is the same as ${\mathcal{M}}$.
By definition, for $\alpha,\beta\in C^{\infty}({\mathcal{L}}_{-\phi})$, there exist $f,g\in C^{\infty}({\mathcal{M}})$ such that $\alpha={\rm e}^{-\operatorname{ad}\phi}\operatorname{pr}^{*}f$ and $\beta={\rm e}^{-\operatorname{ad}\phi}\operatorname{pr}^{*}g$. The derived bracket of $\alpha$ and $\beta$ is equal to
$\big{\{}\big{\{}{\rm e}^{-\operatorname{ad}\phi}\operatorname{pr}^{*}f,{\widetilde{\Theta}}\big{\}},{\rm e}^{-\operatorname{ad}\phi}\operatorname{pr}^{*}g\big{\}}$ $:={\rm e}^{-\operatorname{ad}\phi}\big{\{}\big{\{}\operatorname{pr}^{*}f,{\rm e}^{\operatorname{ad}\phi}{\widetilde{\Theta}}\big{\}},\operatorname{pr}^{*}g\big{\}}.$ (7.2)
@@
$\{\alpha,\beta\}_{{\mathcal{L}}_{-\phi}}$ $:=\operatorname{pr}_{*}{\rm e}^{-\operatorname{ad}\phi}\big{\{}\big{\{}\operatorname{pr}^{*}f,{\rm e}^{\operatorname{ad}\phi}{\widetilde{\Theta}}\big{\}},\operatorname{pr}^{*}g\big{\}}.$ (7.3)
Since it is equivalent to $\operatorname{pr}_{*}\big{\{}\big{\{}{\rm e}^{-\operatorname{ad}\phi}\operatorname{pr}^{*}f,{\widetilde{\Theta}}\big{\}},{\rm e}^{-\operatorname{ad}\phi}\operatorname{pr}^{*}g\big{\}}$, then twisting by $\phi$ gives a graded Poisson bracket on ${\mathcal{L}}_{-\phi}$ induced by ${\widetilde{\Theta}}$ equivalent to a graded Poisson bracket on ${\mathcal{M}}$ induced by ${\rm e}^{\operatorname{ad}\phi}{\widetilde{\Theta}}$. We can prove that {,}$\{-,-\}_{{\mathcal{L}}_{-\phi}}$ is a graded Poisson bracket of degree $-n+1$ similar to the proof for the bracket (7.1).
Thus, in the preceding discussion, we analyze the twisted homological function ${\rm e}^{\operatorname{ad}\phi}{\widetilde{\Theta}}$ on ${\mathcal{M}}$ instead of a homological function on ${\mathcal{L}}_{-\phi}$.
The twisting function $\phi$ is an almost homological function. By (7.2) and (7.3), the graded Poisson bracket is defined by the derived bracket $\big{\{}\big{\{}-,{\widetilde{\Theta}}\big{\}},-\big{\}}$. On the other hand, we obtain
$\operatorname{pr}_{*}{\rm e}^{\operatorname{ad}\phi}{\widetilde{\Theta}}$
$\qquad=\operatorname{pr}_{*}\Big{(}{\widetilde{\Theta}}+\big{\{}{\widetilde{\Theta}},\phi\big{\}}+\frac{1}{2}\big{\{}\big{\{}{\widetilde{\Theta}},\phi\big{\}},\phi\big{\}}+\cdots+\frac{1}{n!}\underbrace{\big{\{}\dots\big{\{}\big{\{}{\widetilde{\Theta}},\phi\big{\}},\phi\big{\}},\dots,\phi\big{\}}}_{n\,\text{brackets}}+\cdots\Big{)}.$ (7.4)
The function $\phi$ is of degree $n$, and from equation (7.1), the third term of $\operatorname{pr}_{*}\,{\rm e}^{\operatorname{ad}\phi}\Theta$ gives
$\{\phi,\phi\}_{{\mathcal{M}}}=-\operatorname{pr}_{*}\big{\{}\big{\{}{\widetilde{\Theta}},\phi\big{\}},\phi\big{\}}.$ (7.5)
If equation (7.5) is zero, $\phi$ is a homological function on ${\mathcal{M}}$ and (,{,}$({\mathcal{M}},\{-,-\}_{{\mathcal{M}}},\phi)$ is a QP-manifold of degree $n-1$. However, equation (7.5) is not necessarily zero.
Inspired by the above analysis and equation (7.4), we define a twisted QP-structure. We consider the following class of generalizations of a QP-manifold which has good properties but it is not a QP.
Definition 7.3.
@@
$\operatorname{pr}_{*}{\rm e}^{\operatorname{ad}\phi}\Theta=0,$ (7.6)
(,{,}$({\mathcal{M}},\{-,-\}_{{\mathcal{M}}},\phi)$ is called a twisted QP-manifold, where $\operatorname{pr}\colon T^{*}[n]{\mathcal{M}}\rightarrow{\mathcal{M}}$ is a natural projection.
A function $\phi$ satisfying the condition (7.6) is also called a a Poisson function or a canonical function. Obviously a QP-manifold is a twisted QP-manifold. A twisted QP-manifold has some applications to physical theories such as the reduction of the AKSZ sigma models and the supergeometric construction of current algebras [3, 28, 29].